Fermi National Laboratory
 

Accelerator Update for May 23, 2006 - May 24, 2006

 

Tuesday May 23

The day shift began with Tevatron (TeV) in controlled access and a Linac RF station (LRF1) down.

An RF expert reported that a Main Injector (MI) RF station (MIRF13) had its power amplifier (PA) fail. Since the repair would require an access, they will wait to make repairs until several tunnel repairs can be lumped together.

Technicians completed their repairs on LRF1 at 8:54 AM. Linac came back online.

Operations established beam to MI at 9:11 AM.

A Booster expert began tuning the machine at 10:04 AM.

Operations established beam to MiniBooNE at 11:17 AM.

At 11:44 AM, MI experts reported that they were having trouble accelerating beam.

A Recycler expert began trying to get beam into the Recycler at 11:48 AM.

A MiniBooNE expert began a target scan study at 1:23 PM.

Cryo system techs began a full cooldown at TeV sector F4 at 1:50 PM.

MI experts successfully accelerated beam to 120 GeV at 3:12 PM.

Recycler succeeded in circulating beam at 4 PM.

Booster studies ended at 4:40 PM.

At 6:07 PM, operators reported that a TeV sector B11 separator vacuum was going bad. Experts were called in to investigate.

Operators had the Pbar rings and transport lines interlocked by 8:05 PM.

At 8:53 PM, Operations reported that the F4 cooldown had stalled and a kautzky valve was stuck. The Cryo Coordinator investigated.

At 9:16 PM, Operations prepared F sector for access.

Cryo techs reported that they had replaced the bad kautzky valve on the F44-2 dipole magnet at 10:57 PM.

Wednesday May 24

The midnight shift began with Operations getting TeV ready to turn back on.

At 12:28 AM, Operations reported a problem with the Transfer Hall safety system. An Interlock tech was called. The tech reported at 2:03 AM that he had found and repaired a bad Keytree connection.

Operations resumed TeV separator conditioning at 2:08 AM.

At 6:44 AM, Booster experts accessed the Booster to add a shunt to ORBMP. Operations turned Booster back on at 7:53 AM.

The Plans for Today

The plans are to allow MI experts to access their tunnels. The Run Coordinator will allow an access tomorrow to fix the P2 vacuum leak.

Machine Reports

Linac

Everything is running normally. LRF station 7 work continues.

Booster

The machine is running well. ORBMP work continues, as does tuning. They hope to have a significant increase of intensity by next week Tuesday.

MI

Experts want to access today to investigate a problem with a transductor at 608 and check other beamline components.

Pbar

The PreVault work should be finished today. They have had the Debuncher extractor kicker on at its nominal value. Techs will replace a bad PA today. Techs will access the tunnel to repair a bad beam valve. CUB will conduct a Pbar LCW pump test today.

TeV

The TeV RF system is up and running. Experts will conduct a Linux sequencer test tomorrow.

Recycler

Experts reported that their turn on was difficult due to MI problems and to a Recycler magnet that had a wrong polarity. They checked their instrumentation and measured their tunes. The electron cooling gas handling leaks were repaired. They are ready for more protons.

Switchyard 120

Experts want to access the P3 line and search for a ground fault on Thursday, 5/25/06.

MiniBooNE

Experts are ready for more beam to conduct target scans.

NuMI

Experts have commissioned their chillers. Electrician work will continue through to Friday. They hope to conduct power supply tests next Tuesday, 5/30/06.

Collider Detector at Fermilab

The work on the southwest shield blocks should be finished this morning. Experts reported that they have a problem with the detector’s central tracker; investigations continue.

D-Zero

They will have the south end of the detector closed today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/24/2006    email Fermilab